OpenStudy (anonymous):

Find the value of the expression |2x-8| x= -2.5; 0; 4;5; 9.5.

OpenStudy (bahrom7893):

just plug in each value instead of x

OpenStudy (anonymous):

Just plug x in for its given values.

OpenStudy (anonymous):

|2x-8| x= -2.5 => -5 |2x-8| x= 0 => -8 |2x-8| x= 4 => 0 |2x-8| x= 5 => 2 |2x-8| x= 9.5 => 11

OpenStudy (bahrom7893):

wait luwcey i think nishnish didn't mean that.

OpenStudy (bahrom7893):

x=-2.5 |2x-8|=|-5-8|=|-13|=13

OpenStudy (anonymous):

oh right! my bad!

OpenStudy (anonymous):

|2x-8| x= -2.5 => l-5l = 5 |2x-8| x= 0 => l-8l =8 |2x-8| x= 4 => 0 |2x-8| x= 5 => 2 |2x-8| x= 9.5 => 11